You are contracting with a wireless Internet service for your home.  The distance from your house to the transmission tower is 35Km an the speed of the link will be 100Mbps. What is the latency for packets of size 2400bytes?
Answer #1

GIven: Distance = 35 Km

Bandwidth = 100 Mbps

Packet size = 2400 bytes.

Latency = Propagation Time + Transmission time + queing time + processing delay

Propagation Time = Distance / Propagation Speed .

Propagation Time = 35000 / (3 * 108 ) (Propagation speed in air = 3 * 108 )

Propagation Time = 0.0001167 =116.7 µs

Transmission Time = Message Size / Bandwidth

Transmission Time = (2400 * 8) / 108

Transmission Time = 192 µs

if we consider negligible queing time and processing delay

Latency = 116.7 + 192 + 0 + 0

Latency = 308.7µs